A) f(0)
B) f(x)=0
f(x)=4x+7
Answer by nyc_function(2741) 
 I will do one section at a time.
For (A), it is telling you replace x with 0 in the given function and solve for
f(0).
f(0) = 4(0) + 7
f(0) = 0 + 7
f(0) = 7
Done!
===================
I will now solve B.
For (B), it is telling you that the entire function equals zero.
So, set 4x + 7 to zero and solve for x.
Let f(x) = 0
f(x) = 4x + 7...This is your function.
0 = 4x + 7
-7 = 4x
-7/4 = x
Done!
|
|
|
| |